OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x))

OpenStudy (anonymous):

f(g(x)) = f(-x +3) = 9(-x+3) -2 = 25 -9x

OpenStudy (anonymous):

Substitute -x+3 in the expression of f(x). Therefore : \[f(g(x))= 9*g(x)-2 = 9*(-x+3)-2\]